An experimental study was conducted by a researcher to determine if a new time slot has an effect on the performance of pupils in mathematics. Fifteen randomly selected learners participated in the study. Toward the end of the investigation, a standardized assessment was conducted. The sample mean was 𝑿̅ = 𝟖𝟓 and the standard deviation s = 3. in the standardization of the test, the mean was 75 and the standard deviation was 10. based on the evidence at hand, is the new timeslot effective?


1
Expert's answer
2022-02-14T16:26:00-0500

Let "X" have a normal distribution with mean "\mu_X" and variance "\sigma_X^2."

Let "Y" have a normal distribution with mean "\mu_Y" and variance "\sigma_Y^2."

If "X" and "Y"are independent, then "Z=X-Y"will follow a normal distribution with mean "\mu_X-\mu_Y" and

variance "\sigma_X^2+\sigma_Y^2."

"\mu_{X-Y}=85-75=10"

"s_{X-Y}=\sqrt{(3)^2+(10)^2}=\sqrt{109}"

The following null and alternative hypotheses need to be tested:

"H_0:\mu=0"

"H_1:\mu\not=0"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, the sample standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05,"

"df=n-1=15-1=14" degrees of freedom, and the critical value for a two-tailed test is "t_{0.025,14}= 2.145"

The rejection region for this two-tailed test is "R=\{t:|t|\gt 2.145\}"

The t-statistic is computed as follows:



"t=\dfrac{\mu_{X-Y}-\mu}{s_{X-Y}/\sqrt{n}}=\dfrac{10-0}{\sqrt{109}\over\sqrt{15}}\approx3.71"


Now, "t=3.71\gt t_{0.025,14}=2.145" thus, we reject the null hypothesis and conclude that the data provides sufficient evidence to show that the new time slot is effective at 5% significance level.